OpenStudy (anonymous):

you inherit $10,000 with the stipulation that for the first year the money must be invested in two stocks paying 6% and 11% annual interest, respectively.how much should be invested at each rate if the total interest earned for the year is to be $700.

OpenStudy (amistre64):

(10000-b)(.11) + b(.06) = 700 10000(.11) -b(.11) +b(.06) = 700 amount@6% = (700 - 10000(.11))/(.06-.11)

OpenStudy (amistre64):

and of course whats leftover is the other amount ....